How do you solve 2/3 = -2x/3 + 3/4 ?
Thank you

Answer by edjones(8007)

2/3 = -2x/3 + 3/4
8=-8x+9 Eliminate the fraction by multiplying each side by 12 (LCM)
-8x=-1
x=1/8
.
Ed